solve by the elimination method 0.1x+0.006y=0.5 and 0.3x-0.2y=3.4
Answer by Alan3354(69443)

0.1x+0.006y=0.5 x30 --> 3x + 0.18y = 15
0.3x -0.2y =3.4 x10 --> 3x - 2y = 34
=========================================
3x + 0.18y = 15
3x - 2y = 34
-------------------- Subtract
2.18y = -19
y = -950/109
--------
Sub for y to find x